According to a​ magazine, as of January 2018, 6% of chief executive officers were women. Complete parts A through D based on a random sample of 12 CEO's.

a. What is the probability that one corporate officer was​ female?

b. What is the probability that fewer than four corporate officers were​ female?

c. What is the probability that more than two corporate officers were​ female?

d. What are the mean and standard deviation for this​ distribution?
Answer #1

Given probability of female chief executive officers are p -0.06 Sample size n=12 Xhas binomial distribution with the parametc) Probability of more than 2 officers were female is P(x>2) 1-P(x2) %3D1-{Р(х-0)-Р9х-1)+P(х-2)} = 1-{0.4759+0.3645+0.1280 =